OpenStudy (anonymous):

3. How many carbon atoms are contained in 0.5 mol glucose, C6H12O6 ?

OpenStudy (abb0t):

Use avogadro's number: \(6.022 \times 10^{23} \large \frac{atoms}{mol}\)

OpenStudy (nirmalnema):

C6H12O6 means 1 mol of glucose...and it contains.......6 *6.022*10^23 carbon atoms.. therefore...0.5 mol has 6/1*.0.5*6.022*10^23 carbon atoms which = 18.066*10^23 carbon atoms

OpenStudy (abb0t):

I cannot understand anything you wrote up there. Try re-writing it in LaTex.